1.Use the principle of homogeneity, check the correctness of the following equation

β„Ž=2π‘‡πΆπ‘œπ‘ π›½/π‘Ÿπ‘‘π‘”

where Ξ² is the angle of contact, d is the density of a liquid, r is the radius of the tube, g is acceleration due to gravity, h is the height of the liquid and T is the surface tension.

2. If the escape velocity depends upon;

(i) Acceleration due to gravity of the planet

(ii) The radius of the planet

Establish the relationship between them by dimensional analysis.

[Hint, 𝑣𝑒𝛼 π‘”π‘Žπ‘…π‘]



1
Expert's answer
2021-03-10T17:22:18-0500

1)


"[L]=\frac{[MT^{-2}]}{[ML^{-3}][L][LT^{-2}]}\\ [L]=[L]"

So, the relation is correct.

2)


"v=g^aR^b\\ [v]=[g]^a[R]^b\\LT^{-1}=(LT^{-2})^a(L)^b\\a=0.5,b=0.5\\
v=k\sqrt{Rg}"